OpenStudy (anonymous):

Am I right? [sqrt(10)(cos(40)+i sin(40)]^6 I need to put it in a+bi format but I just keep getting 1000??

OpenStudy (anonymous):

From DeMoivre's theorem, you get \[\left(\sqrt{10}(\cos40+i\sin40)\right)^6=10^3(\cos240+i\sin240)\] Check your sine/cosine again.

OpenStudy (anonymous):

I got, -500-500isaqrt(3)

OpenStudy (anonymous):

That's correct.

OpenStudy (anonymous):

that's the answer? It doesn't have to be positive then?

OpenStudy (anonymous):

That's right, and no, if you recall the unit circle, an angle of \(240^\circ\) will land you in the third quadrant, which contains points with negative \(x\) and \(y\) components.

OpenStudy (anonymous):

Thank you, do you think you can also help me in another question? I'll open a new thread so I can give you another medal.

OpenStudy (anonymous):

Sure

OpenStudy (anonymous):

yw